# AL-Bench: A Benchmark for Automatic Logging

### v1

@@ 4.3. RQ2: How well can the predicted log statements be compiled successfully?
To evaluate tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances of compile failure. When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking. To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
@@ 4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
Table 6. The Semantic Similarity of Logs Printed by the Predicted Log Statements Across Methods.
@@
Answer to RQ3. The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
@@ 5. Threats to Vadility
  • External Validity: In this work we evaluate the tools with 10 widely recognized Java projects in different domains. The results of each tools are examined in different requirements of logging practices. The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.

### v2

@@ 4.3. RQ2: How well can the predicted log statements be compiled successfully?
To evaluate the tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances of compile failure. When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking. To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
@@ 4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
Table 6. The semantic similarity between logs printed by the predicted log statements and the ground truth.
@@
Answer to RQ3. The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting a higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
@@ 5. Threats to Vadility
  • External Validity: In this work, we evaluate the tools with 10 widely recognized Java projects in different domains. The results of each tool are examined in different requirements of logging practices. The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.